# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v2

@@ Appendix D Algorithm of CompFlow built on Soft-Actor-Critic
We will use $\nu:=\begin{bmatrix}s\\ a\end{bmatrix}$ to denote the state action pair. According to [57], for an estimator $\hat{s}^{\prime}$ generated by a linear path flow, the squared error admits the bound
$\|s^{\prime}-\hat{s}^{\prime}\|^{2}\;<\;\mathrm{Tr}(\Sigma_{s^{\prime}})-\mathrm{Tr}\!\bigl{(}\Sigma_{s^{\prime}\nu}\Sigma_{\nu}^{-1}\Sigma_{\nu s^{\prime}}\bigr{)}\;=\;\mathrm{Var}(s^{\prime}\mid\nu),$
where $\Sigma_{s^{\prime}}$, $\Sigma_{\nu}$, and $\Sigma_{s^{\prime}\nu}$ are the covariance and cross-covariance matrices of $s^{\prime}$ and $\nu$.
Using the $N_{\text{tar}}$ target samples to train the direct flow, we have
$Err_{\text{direct}}=\mathrm{Var}(s^{\prime}\mid\nu)_{\text{tar}}+O\!\Bigl{(}\tfrac{\sigma_{\text{tar}}^{2}}{\sqrt{N_{\text{tar}}}}\Bigr{)},\quad\sigma_{\text{tar}}^{2}:=\mathrm{Tr}(\Sigma_{s^{\prime}}^{\text{tar}}).$
The composite flow first trains on the $N_{\text{src}}$ source samples and then adapts to the target distribution. Standard covariance concentration therefore gives
$Err_{\text{cmpst}}\leq\mathrm{Var}(s^{\prime}\mid\nu)_{\text{src}}+O\!\Bigl{(}\tfrac{\sigma_{\text{src}}^{2}}{\sqrt{N_{\text{src}}}}\Bigr{)},\quad\sigma_{\text{src}}^{2}:=\mathrm{Tr}(\Sigma_{s^{\prime}}^{\text{src}}).$
For $Err_{\text{cmpst}}<Err_{\text{direct}}$, a sufficient condition is
$\mathrm{Var}(s^{\prime}\mid\nu)_{\text{src}}-\mathrm{Var}(s^{\prime}\mid\nu)_{\text{tar}}\;<\;O\!\Bigl{(}\tfrac{\sigma_{\text{tar}}^{2}}{\sqrt{N_{\text{tar}}}}-\tfrac{\sigma_{\text{src}}^{2}}{\sqrt{N_{\text{src}}}}\Bigr{)}.$ (8)
Now we related the variance gap between source and target transitions to total variation distance between source and target transitions. Consider $s^{\prime}\in\mathbb{R}^{d}$ be a random vector. Let $P$, $Q$ be two probability distributions on $\mathbb{R}^{d}$, and recall that $|s^{\prime}|\leq\zeta$.
The variance difference is as follows and we bound it clearly by triangle inequality:
$|\text{Var}_{P}(s^{\prime})-\text{Var}_{Q}(s^{\prime})|$ $=|\mathbb{E}_{P}[s^{\prime 2}]-(\mathbb{E}_{P}[s^{\prime}])^{2}-(\mathbb{E}_{Q}[s^{\prime 2}]-(\mathbb{E}_{Q}[s^{\prime}])^{2})|$ (9)
$=|(\mathbb{E}_{P}[s^{\prime 2}]-\mathbb{E}_{Q}[s^{\prime 2}])-[(\mathbb{E}_{P}[s^{\prime}])^{2}-(\mathbb{E}_{Q}[s^{\prime}])^{2}]|$ (10)
$\leq|\mathbb{E}_{P}[s^{\prime 2}]-\mathbb{E}_{Q}[s^{\prime 2}]|+|(\mathbb{E}_{P}[s^{\prime}])^{2}-(\mathbb{E}_{Q}[s^{\prime}])^{2}|.$ (11)
For the first term:
@@
since $|s^{\prime}|^{2}\leq\zeta^{2}$ (bounded state space) and the absolute difference is directly bounded by total variation norm $\|P-Q\|_{1}$.
For the second term, explicitly factor it:
$|(\mathbb{E}_{P}[s^{\prime}])^{2}-(\mathbb{E}_{Q}[s^{\prime}])^{2}|=|\mathbb{E}_{P}[s^{\prime}]-\mathbb{E}_{Q}[s^{\prime}]|\cdot|\mathbb{E}_{P}[s^{\prime}]+\mathbb{E}_{Q}[s^{\prime}]|.$
Each expectation is bounded by $\zeta$ since $|s^{\prime}|\leq\zeta$:
$|(\mathbb{E}_{P}[s^{\prime}])^{2}-(\mathbb{E}_{Q}[s^{\prime}])^{2}|\leq|\mathbb{E}_{P}[s^{\prime}]-\mathbb{E}_{Q}[s^{\prime}]|\cdot 2\zeta.$
Moreover, explicitly bounding the expectation difference by total variation distance (standard property of expectations w.r.t. total variation):
[…]
Thus rigorously, the second term is bounded as:
$|(\mathbb{E}_{P}[s^{\prime}])^{2}-(\mathbb{E}_{Q}[s^{\prime}])^{2}|\leq 2\zeta^{2}\|P-Q\|_{1}.$
Combining both terms explicitly yields:
$|\text{Var}_{P}(s^{\prime})-\text{Var}_{Q}(s^{\prime})|\leq 3\zeta^{2}\|P-Q\|_{1}$
Applying this with $P=p_{\text{src}}(\cdot\mid\nu)$ and $Q=p_{\text{tar}}(\cdot\mid\nu)$ and taking the supremum over $\nu$ gives
$\mathrm{Var}(s^{\prime}\mid\nu)_{\text{src}}-\mathrm{Var}(s^{\prime}\mid\nu)_{\text{tar}}\;\leq\;3\zeta^{2}\,\sup_{\nu}D_{\mathrm{TV}}\left(p_{\mathrm{src}}(\cdot\mid\nu),\;p_{\mathrm{tar}}(\cdot\mid\nu)\right).$
We thus get a sufficient condition for the composite flow to outperform the direct flow is
$3\zeta^{2}\,\sup_{\nu}D_{\mathrm{TV}}\left(p_{\mathrm{src}}(\cdot\mid\nu),\;p_{\mathrm{tar}}(\cdot\mid\nu)\right)\;<\;O\!\Bigl{(}\tfrac{\sigma_{\text{tar}}^{2}}{\sqrt{N_{\text{tar}}}}-\tfrac{\sigma_{\text{src}}^{2}}{\sqrt{N_{\text{src}}}}\Bigr{)},$
or, equivalently,
$D_{TV}(p_{\text{src}},p_{\text{tar}})=O\!\Bigl{(}\tfrac{1}{\zeta^{2}}(\tfrac{\operatorname{Tr}(\Sigma_{s^{\prime}}^{\text{tar}})}{\sqrt{N_{\text{tar}}}}-\tfrac{\operatorname{Tr}(\Sigma_{s^{\prime}}^{\text{src}})}{\sqrt{N_{\text{src}}}})\Bigr{)},$
which is precisely the condition stated in the theorem. ∎
Appendix G Proof of Theorem 3.3
To prove theorem  3.3, we first prove the following lemma which bounds the Wasserstein-2 distance estimation error of the optimal transport flow matching model by the square root of its mean square error.
Lemma G.1 (Wasserstein-2 Distance Estimation Bound).
Let $p_{\mathrm{src}},p_{\mathrm{tar}}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ be source and target distributions that also satisfies assumptions (A1) and (A2) in Proposition 3.2 (see Appendix H). Let $\hat{T}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a learned transport map obtained by flow matching, and define the pushforward distribution $\hat{p}_{\mathrm{tar}}:=\hat{T}_{\#}p_{\mathrm{src}}$. Then the error in estimating the Wasserstein-2 distance between the source and target distributions using $\hat{T}$ is bounded by the square root of the mean squared error between $\hat{T}(x_{0})$ and the true matched target sample $s$, i.e.,
$\left|W_{2}(p_{\mathrm{src}},p_{\mathrm{tar}})-W_{2}(p_{\mathrm{src}},\hat{p}_{\mathrm{tar}})\right|\leq\sqrt{\mathrm{Err}_{\mathrm{MSE}}}.$
@@
Note that $W_{2}(\cdot,\cdot)$ is a metric, so by the triangle inequality, we have:
$|W_{2}(p_{\text{src}},p_{\text{tar}})-W_{2}(p_{\text{src}},\hat{p}_{\text{tar}})|\leq W_{2}(p_{\text{tar}},\hat{p}_{\text{tar}}).$
Because assumptions (A1) and (A2) in Proposition 3.2 are satisfied, by Brenier theorem there exists a deterministic optimal transport map $T$. We denote a sample from the source distribution $x_{0}$ and the matching target sample induced by the true optimal transport map $T$ as $s:=T(x_{0})$, and the matching target sample induced by the learned optimal transport map $\hat{T}$ as $\hat{s}:=\hat{T}(x_{0})$.
We consider $(x_{0},s)\sim q$, where $q$ is a coupling between the real source and target distributions used in training the flow matching model, with marginals $x_{0}\sim p_{\text{src}}$ and $s\sim p_{\text{tar}}$. Let $\Pi(\hat{p}_{\rm tar},p_{\rm tar})$ denote the set of all couplings between $\hat{p}_{\rm tar}$ and $p_{\rm tar}$. By the definition of $W_{2}$, we have:
$W_{2}^{2}(p_{\text{tar}},\hat{p}_{\text{tar}})=\inf_{q^{\prime}\in\Pi(p_{\text{tar}},\hat{p}_{\text{tar}})}\mathbb{E}_{(\hat{s},s)\sim q^{\prime}}\|\hat{s}-s\|^{2}\leq\mathbb{E}_{(x_{0},s)\sim q}\|\hat{T}(x_{0})-s\|^{2}$
So:
$W_{2}(p_{\text{tar}},\hat{p}_{\text{tar}})\leq\sqrt{\mathbb{E}_{(x_{0},s)\sim q}\|\hat{T}(x_{0})-s\|^{2}}=\sqrt{\text{Err}_{\text{MSE}}}$
@@
See 3.3
@@
We list the full description of assumptions made in the theorem statement:
1. Lipschitz reward. For every $a\in\mathcal{A}$ the map $s\mapsto r(s,a)$ is $L_{r}$-Lipschitz:
$|r(s,a)-r(s^{\prime},a)|\;\leq\;L_{r}\,d(s,s^{\prime})\quad\forall s,s^{\prime},a.$
2. Lipschitz dynamics. Both kernels are $L_{p}$-Lipschitz in Wasserstein distance:
$W_{1}\!\bigl{(}p_{m}(\cdot\mid s,a),p_{m}(\cdot\mid s^{\prime},a)\bigr{)}\;\leq\;L_{p}\,d(s,s^{\prime})\quad\forall s,s^{\prime},a,\;m\in\{\text{src},\text{tar}\}.$
3. Contraction margin. $\gamma L_{p}<1$.
Let’s first study the performance of a policy $\pi$ trained with the source MDP dataset on the target MDP $\eta_{\mathcal{M}_{\text{tar}}}\!\bigl{(}\pi\bigr{)}$. Denote $\Delta_{W_{1}}\;:=\;\sup_{s,a}W_{1}\!\bigl{(}p_{\text{src}}(\cdot|s,a),\;p_{\text{tar}}(\cdot|s,a)\bigr{)}$.
Given a fixed policy $\pi$, we cite Theorem 3.1 in  [40] for the following bound:
$\left\|V^{\pi}_{\text{tar}}-V^{\pi}_{\text{src}}\right\|_{\infty}$ $\leq 2L_{r}\Delta_{W_{1}}(1+\gamma)\sum_{i=0}^{\infty}\gamma^{i}\sum_{j=0}^{i}(L_{P})^{j}$
@@
$\bigl{|}\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\mathcal{M}_{\text{src}}}(\pi)\bigr{|}$ $=\Bigl{|}\int_{\mathcal{S}}\bigl{(}V^{\pi}_{\mathrm{tar}}(s)-V^{\pi}_{\mathrm{src}}(s)\bigr{)}\,\mu(ds)\Bigr{|}$
$\leq\int_{\mathcal{S}}\bigl{|}V^{\pi}_{\mathrm{tar}}(s)-V^{\pi}_{\mathrm{src}}(s)\bigr{|}\,\mu(ds).$
$\leq\left\|V^{\pi}_{\text{tar}}-V^{\pi}_{\text{src}}\right\|_{\infty}$
@@
Moreover, the bounded state space assumption in Theorem  3.1 implies bounded second moment for both $p_{\text{src}}$ and $p_{\text{tar}}$, so we have $W_{1}(p_{\text{src}},p_{\text{tar}})\leq W_{2}(p_{\text{src}},p_{\text{tar}})\ \forall p_{\text{src}},p_{\text{tar}}$. Hence, we obtain that
$|\eta_{\mathcal{M}_{\text{tar}}}(\pi)-\eta_{\mathcal{M}_{\text{src}}}(\pi)|$ $\;\leq\;\frac{2L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\Delta_{W_{1}}$ (13)

### v3

@@ Appendix E Algorithm of CompFlow built on Soft-Actor-Critic
We consider linear-path flow matching (FM) trained with squared loss. A single training example is generated by: $t\sim\mathrm{Unif}[0,1]$; $X_{0}\sim p_{0}$ (the start distribution), independently of $t$; $X_{1}\sim p_{\rm on}$, independently of $(t,X_{0})$; The interpolation $X_{t}:=(1-t)X_{0}+tX_{1}$ and the label $Y:=X_{1}-X_{0}\in\mathbb{R}^{d}$.
Let $\mathcal{H}$ be a hypothesis class of (measurable) vector fields $v:\mathbb{R}^{d}\times[0,1]\to\mathbb{R}^{d}$ and define the population risk
$R(v):=\mathbb{E}\!\left[\,\left\lVert v(X_{t},t)-Y\right\rVert_{2}^{2}\,\right],$ (10)
where the expectation is over the sampling mechanism above. Let
$v^{\star}(x,t):=\mathbb{E}\!\left[\,Y\,\middle|\,X_{t}=x,\,t\,\right]$ (11)
denote the Bayes (population) minimizer. Given $n$ i.i.d. samples, let $\hat{v}$ be any empirical risk minimizer over $\mathcal{H}$ for the squared loss.
Assumptions. A1 (Uniform boundedness) There exists $B\in(0,\infty)$ such that $\sup_{v\in\mathcal{H}}\sup_{x\in\mathbb{R}^{d},\,t\in[0,1]}\left\lVert v(x,t)\right\rVert_{2}\leq B$. This ensures integrability and controls the Lipschitz constants used below.
FM as regression; Bayes predictor and excess-risk identity.
By definition,
$R(v)\;=\;\mathbb{E}\,\left\lVert v(X_{t},t)-Y\right\rVert_{2}^{2},\quad Y=X_{1}-X_{0}.$
For any measurable $v$, the Bayes (population) minimizer for the squared loss is the conditional mean
$v^{\star}(x,t)\;=\;\mathbb{E}\!\left[\,Y\,\middle|\,X_{t}=x,\,t\,\right].$
A standard identity for squared loss (we derive it fully) is
$R(v)-R(v^{\star})\;=\;\mathbb{E}\,\left\lVert\,v(X_{t},t)-v^{\star}(X_{t},t)\,\right\rVert_{2}^{2}.$ (12)
Derivation of (12). Expand and add/subtract $v^{\star}(X_{t},t)$:
$R(v)$ $=\mathbb{E}\,\left\lVert\,v(X_{t},t)-v^{\star}(X_{t},t)+v^{\star}(X_{t},t)-Y\,\right\rVert_{2}^{2}$
$=\mathbb{E}\,\left\lVert v-v^{\star}\right\rVert_{2}^{2}\;+\;2\,\mathbb{E}\,\langle v-v^{\star},\,v^{\star}-Y\rangle\;+\;\mathbb{E}\,\left\lVert v^{\star}-Y\right\rVert_{2}^{2},$
where all functions are evaluated at $(X_{t},t)$ but notationally suppressed for readability. Condition on $(X_{t},t)$: by definition, $\mathbb{E}[Y\mid X_{t},t]=v^{\star}(X_{t},t)$, hence $\mathbb{E}[\,v^{\star}-Y\mid X_{t},t\,]=0$ and the cross term vanishes after taking expectations. Therefore
$R(v)=\mathbb{E}\,\left\lVert v-v^{\star}\right\rVert_{2}^{2}+R(v^{\star}),$
which rearranges to (12).
In Lemma G.3, we prove that there exists an absolute constant $c>0$ such that, for all $\delta\in(0,1)$, with probability at least $1-\delta$ (over the sample draw),
$R(\hat{v})\;\leq\;R(v^{\star})\;+\;c\,\Big(B+\sqrt{\mathbb{E}\,\left\lVert Y\right\rVert_{2}^{2}}\Big)\,\Gamma_{n,\delta},\qquad\Gamma_{n,\delta}:=\mathfrak{R}_{n}(\mathcal{H})\;+\;\sqrt{\tfrac{\log(1/\delta)}{n}},$ (13)
where $\mathfrak{R}_{n}(\mathcal{H})$ is the (data-independent) Rademacher complexity of $\mathcal{H}$. In short: symmetrization ++ vector contraction ++ Lipschitzness of the squared loss on the $B$-ball yields (13).)
Subtract $R(v^{\star})$ on both sides of  (13) and apply (12) with $v=\hat{v}$:
$\mathbb{E}\,\left\lVert\hat{v}-v^{\star}\right\rVert_{2}^{2}\;\leq\;c\,\Big(B+\sqrt{\mathbb{E}\,\left\lVert Y\right\rVert_{2}^{2}}\Big)\,\Gamma_{n,\delta}.$ (14)
Identify $\mathbb{E}\,\left\lVert Y\right\rVert_{2}^{2}$ as $D_{2}(p_{0},p_{\rm on})$ and linearize the square root.
By independence of $X_{0}\!\sim p_{0}$ and $X_{1}\!\sim p_{\rm on}$,
$\mathbb{E}\,\left\lVert Y\right\rVert_{2}^{2}\;=\;\mathbb{E}\,\left\lVert X_{1}-X_{0}\right\rVert_{2}^{2}\;=:\;D_{2}(p_{0},p_{\rm on}).$ (15)
By applying Lemma G.4, we have
$\sqrt{\mathbb{E}\,\left\lVert Y\right\rVert_{2}^{2}}\leq\sqrt{\operatorname{Tr}(\Sigma_{\rm on})+C_{\rm TR}}+W_{2}(p_{0},p_{\rm on}).$ (16)
Therefore, we obtain
$\mathbb{E}\,\left\lVert\hat{v}-v^{\star}\right\rVert_{2}^{2}$ $\leq c\,\Big(B+\sqrt{\operatorname{Tr}(\Sigma_{\rm on})+C_{\rm TR}}+W_{2}(p_{0},p_{\rm on})\Big)\,\Gamma_{n,\delta}$ (17)
This gives the following master bound (18) with $C_{0}:=c(B+\sqrt{\mathrm{Tr}(\Sigma_{\rm on})+C_{\rm TR}}$ and $C_{1}:=c$.
Master bound. Under A1, for any start $p_{0}$ (with fixed target $p_{\rm on}$), with probability at least $1-\delta$,
$\mathbb{E}\,\left\lVert\hat{v}-v^{\star}\right\rVert_{2}^{2}\;\leq\;\underbrace{C_{0}\,\Gamma_{n,\delta}}_{\text{$p_{0}$-independent}}\;+\;\underbrace{C_{1}\,W_{2}(p_{0},p_{\rm on})\,\Gamma_{n,\delta}}_{\text{depends on the start }p_{0}},\qquad C_{0}:=c(B+\sqrt{\mathrm{Tr}(\Sigma_{\rm on})+C_{\rm TR}}),\;\;C_{1}:=c.$ (18)
Composite vs. direct. Consider two trainings sharing the same online $p_{\rm on}$ and class $\mathcal{H}$:
(Composite step) $\quad p_{0}\;=\;\hat{p}_{\rm off}(\cdot\mid s,a),$ (19)
(Direct step) $\quad p_{0}\;=\;p_{G}\;=\;\mathcal{N}(0,I_{d}).$ (20)
Applying (18) to (19) and (20) yields
$\mathbb{E}\,\left\lVert\hat{v}_{\mathrm{comp}}-v^{\star}_{\mathrm{comp}}\right\rVert_{2}^{2}$ $\;\leq\;C_{0}\,\Gamma_{n,\delta}+C_{1}\,W_{2}\!\big(\hat{p}_{\rm off},p_{\rm on}\big)\,\Gamma_{n,\delta},$ (21)
$\mathbb{E}\,\left\lVert\hat{v}_{G}-v^{\star}_{G}\right\rVert_{2}^{2}$ $\;\leq\;C_{0}\,\Gamma_{n,\delta}+C_{1}\,W_{2}\!\big(p_{G},p_{\rm on}\big)\,\Gamma_{n,\delta}.$ (22)
When is the composite bound strictly tighter?
Applying (18) to $p_{0}=\hat{p}_{\rm off}$ (composite) and $p_{0}=p_{G}$ (direct) gives (21) and (22). Since $C_{0}$ and $\Gamma_{n,\delta}$ are identical across the two trainings (they depend on $(B,c,\Sigma_{\rm on},C_{\rm TR})$, $n$, $\delta$, and $\mathcal{H}$, but not on $p_{0}$), the only difference in the right-hand sides is the factor $W_{2}(\cdot,p_{\rm on})$.
Therefore the composite bound (21) is strictly smaller than the direct bound (22) if and only if
$W_{2}\!\big(\hat{p}_{\rm off},p_{\rm on}\big)\;<\;W_{2}\!\big(p_{G},p_{\rm on}\big).$ (23)
Remark G.2 (Optional $D_{2}$ strengthening).
One can analogously replace $W_{2}(p_{0},p_{\rm on})$ by $D_{2}(p_{0},p_{\rm on})$ when bounding 14, using the standard inequality trick $\sqrt{x}\leq\tfrac{1}{2}(x+1)$ for all $x\geq 0$. We use $W_{2}$ metric because it’s more commonly adopted in the literature, but $D_{2}$ leads to a tighter bound since we no longer needs the constant $C_{\rm TR}$. We do not expand the $D_{2}$ case here, as the argument mirrors the $W_{2}$ case.
Lemma G.3 (Generalization gap for squared loss under uniform boundedness).
Assume A1. Let $\ell_{v}(z):=\left\lVert v(x,t)-y\right\rVert_{2}^{2}$ for $z=(x,t,y)$ generated as in the theorem. Then there exists a positive constant $c>0$ such that, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over an i.i.d. sample of size $n$,
$R(\hat{v})\;\leq\;R(v^{\star})\;+\;c\,\Big(B+\sqrt{\mathbb{E}\,\left\lVert Y\right\rVert_{2}^{2}}\Big)\,\Gamma_{n,\delta},\qquad\Gamma_{n,\delta}:=\mathfrak{R}_{n}(\mathcal{H})+\sqrt{\tfrac{\log(1/\delta)}{n}}.$
Proof. We sketch the standard steps and make constants explicit where needed:
(i) Symmetrization. Let $\hat{R}(v)$ denote the empirical risk. For ERM $\hat{v}$,
@@
up to negligible terms; this is standard (e.g., by a one-sided symmetrization plus a chaining step). Thus it suffices to bound $\sup_{v\in\mathcal{H}}(R(v)-\hat{R}(v))$.
(ii) Lipschitz envelope via contraction. For fixed $(x,t,y)$ with $\left\lVert v(x,t)\right\rVert\leq B$ and any $u\in\mathbb{R}^{d}$ with $\left\lVert u\right\rVert\leq B$,
$\Big|\,\left\lVert u-y\right\rVert^{2}-\left\lVert v(x,t)-y\right\rVert^{2}\,\Big|=\big|\,\langle u-v(x,t),\,(u+v(x,t)-2y)\rangle\,\big|\leq 2\,(B+\left\lVert y\right\rVert)\,\left\lVert u-v(x,t)\right\rVert.$
Thus, as a function of $u$, the map $u\mapsto\left\lVert u-y\right\rVert^{2}$ is $(2(B+\left\lVert y\right\rVert))$-Lipschitz on the $B$-ball. Applying the vector contraction inequality to the class $\{v(\cdot,\cdot)\in\mathcal{H}\}$ gives
$\mathbb{E}\,\sup_{v\in\mathcal{H}}\bigl(R(v)-\hat{R}(v)\bigr)\;\lesssim\;\Big(\,\mathbb{E}\,[\,B+\left\lVert Y\right\rVert\,]\,\Big)\,\mathfrak{R}_{n}(\mathcal{H}).$
By Cauchy–Schwarz and Jensen, $\mathbb{E}\left\lVert Y\right\rVert\leq\sqrt{\mathbb{E}\left\lVert Y\right\rVert^{2}}$.
(iii) High-probability upgrade. A standard bounded differences (or Bernstein-type) argument for Lipschitz, sub-quadratic losses upgrades the expected sup-gap to a high-probability bound, adding the usual $\sqrt{\log(1/\delta)/n}$ term. Collecting constants and using $\mathbb{E}\left\lVert Y\right\rVert\leq\sqrt{\mathbb{E}\left\lVert Y\right\rVert^{2}}$ yields the claimed form with some absolute $c>0$:
$\sup_{v\in\mathcal{H}}\bigl(R(v)-\hat{R}(v)\bigr)\;\leq\;c\,\Big(B+\sqrt{\mathbb{E}\left\lVert Y\right\rVert^{2}}\Big)\,\Gamma_{n,\delta}.$
Since $\hat{v}$ is an ERM, $\hat{R}(\hat{v})\leq\hat{R}(v^{\star})$, and therefore $R(\hat{v})-R(v^{\star})\leq 2\sup_{v}(R(v)-\hat{R}(v))$ in the final step, absorbing the factor 2 into $c$. $\square$
Lemma G.4.
Let $\mu_{\cdot}$ and $\Sigma_{\cdot}$ denote mean and covariance, respectively. We show that if $\mathrm{Tr}(\Sigma_{0})\leq C_{\rm TR}$ for some $C_{\rm TR}>0$, we have:
$\sqrt{D_{2}(p_{0},p_{1})}\leq\sqrt{\operatorname{Tr}(\Sigma_{1})+C_{\rm TR}}+W_{2}(p_{0},p_{1})$
@@
Let $X_{0}\sim p_{0}$ and $X_{1}\sim p_{1}$ be independent with means $\mu_{0},\mu_{1}$ and covariances $\Sigma_{0},\Sigma_{1}$. We compute
$D_{2}(p_{0},p_{1})$ $=\mathbb{E}\,\left\lVert X_{1}-X_{0}\right\rVert_{2}^{2}=\mathbb{E}\,\left\lVert X_{1}\right\rVert_{2}^{2}+\mathbb{E}\,\left\lVert X_{0}\right\rVert_{2}^{2}-2\,\mathbb{E}\,\langle X_{1},X_{0}\rangle.$
By independence, $\mathbb{E}\,\langle X_{1},X_{0}\rangle=\langle\mathbb{E}X_{1},\mathbb{E}X_{0}\rangle=\langle\mu_{1},\mu_{0}\rangle$. Moreover, for any random vector $Z$ with mean $\mu$ and covariance $\Sigma$, $\mathbb{E}\,\left\lVert Z\right\rVert_{2}^{2}=\left\lVert\mu\right\rVert_{2}^{2}+\mathrm{Tr}(\Sigma)$. Thus
$D_{2}(p_{0},p_{1})$ $=\big(\left\lVert\mu_{1}\right\rVert_{2}^{2}+\mathrm{Tr}(\Sigma_{1})\big)+\big(\left\lVert\mu_{0}\right\rVert_{2}^{2}+\mathrm{Tr}(\Sigma_{0})\big)-2\,\langle\mu_{1},\mu_{0}\rangle$
$=\left\lVert\mu_{1}-\mu_{0}\right\rVert_{2}^{2}+\mathrm{Tr}(\Sigma_{1}+\Sigma_{0}),$
Let $\Gamma(p_{0},p_{1})$ be all couplings of $p_{0},p_{1}$. For any $\pi\in\Gamma(p_{0},p_{1})$ with $(X_{0},X_{1})\sim\pi$, we have
$\mathbb{E}_{\pi}\|X_{1}-X_{0}\|_{2}^{2}\;\geq\;\big\|\mathbb{E}_{\pi}[X_{1}-X_{0}]\big\|_{2}^{2}\;=\;\|\mu_{1}-\mu_{0}\|_{2}^{2},$
by Jensen’s inequality since $z\mapsto\|z\|_{2}^{2}$ is convex. Taking the infimum over all couplings gives
$W_{2}^{2}(p_{0},p_{1})=\inf_{\pi\in\Gamma(p_{0},p_{1})}\mathbb{E}_{\pi}\|X_{1}-X_{0}\|_{2}^{2}\;\geq\;\|\mu_{1}-\mu_{0}\|_{2}^{2}.$
Therefore, we have $\|\mu_{1}-\mu_{0}\|_{2}^{2}\leq W_{2}^{2}(p_{0},p_{1})$. Hence,
$\sqrt{D_{2}(p_{0},p_{1})}$ $=\sqrt{\left\lVert\mu_{1}-\mu_{0}\right\rVert_{2}^{2}+\mathrm{Tr}(\Sigma_{1}+\Sigma_{0})}$
$\leq\sqrt{\mathrm{Tr}(\Sigma_{0}+\Sigma_{1})}+\sqrt{\left\lVert\mu_{1}-\mu_{0}\right\rVert_{2}^{2}}$
$\leq\sqrt{\mathrm{Tr}(\Sigma_{1})+C_{\rm TR}}+W_{2}(p_{0},p_{1})$
@@
The assumption that $\operatorname{Tr}(\Sigma_{0})\leq C_{\rm TR}$ makes sense because we assume bounded state space for all settings considered in our paper.
Appendix H Proof of Theorem 3.4
See 3.4
@@
Let’s first list the assumptions:
R1: Lipschitz reward. For every $a\in\mathcal{A}$, $s\mapsto r(s,a)$ is $L_{r}$-Lipschitz.
R2: Lipschitz dynamics. For $m\in\{\rm off,\rm on\}$, $W_{1}\!\bigl(p_{m}(\cdot\mid s,a),p_{m}(\cdot\mid s^{\prime},a)\bigr)\leq L_{p}\,d(s,s^{\prime})$ for all $s,s^{\prime},a$.
R3: Contraction. $\gamma L_{p}<1$.
We proceed in four steps: (A) a policy-dependent model-gap bound; (B) a uniform high-probability W2-gap estimation bound for the rectified flow ($\alpha$); (C) learning-gap bounds for $\pi_{\mathrm{bc}}$ and $\hat{\pi}$ ($\beta$); (D) assembly via the standard three-term decomposition.
Step A: Policy-dependent model-gap bound.
For any fixed policy $\pi$, we cite Theorem 3.1 of [neufeld2023bounding] for the following bound. Under R1R3, we have
$\left\|V^{\pi}_{\rm on}-V^{\pi}_{\rm off}\right\|_{\infty}$ $\leq 2L_{r}\Delta_{W_{1}}(1+\gamma)\sum_{i=0}^{\infty}\gamma^{i}\sum_{j=0}^{i}(L_{P})^{j}$
@@
$\bigl|\eta_{\mathcal{M}_{\rm on}}(\pi)-\eta_{\mathcal{M}_{\rm off}}(\pi)\bigr|$ $=\Bigl|\int_{\mathcal{S}}\bigl(V^{\pi}_{\rm on}(s)-V^{\pi}_{\rm off}(s)\bigr)\,\mu(ds)\Bigr|$
$\leq\int_{\mathcal{S}}\bigl|V^{\pi}_{\rm on}(s)-V^{\pi}_{\rm off}(s)\bigr|\,\mu(ds).$
$\leq\left\|V^{\pi}_{\rm on}-V^{\pi}_{\rm off}\right\|_{\infty}$
@@
Moreover, the bounded state space assumption in Theorem  3.1 implies bounded second moment for both $p_{\rm off}$ and $p_{\rm on}$, so we have $W_{1}(p_{\rm off},p_{\rm on})\leq W_{2}(p_{\rm off},p_{\rm on})\ \forall p_{\rm off},p_{\rm on}$. Hence, we obtain that
$|\eta_{\mathcal{M}_{\rm on}}(\pi)-\eta_{\mathcal{M}_{\rm off}}(\pi)|$ $\;\leq\;\frac{2L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\Delta_{W_{1}}$ (25)